integrate (1-2x)/(45+4x-x^2) dx without using trig substitution


1
Expert's answer
2021-06-04T05:07:16-0400

"x^2-4x-45=0"

"x=\frac{4\pm \sqrt{16+180}}{2}"

"x_1=-5,x_2=9"

"x^2-4x-45=(x+5)(x-9)"

"\frac{1-2x}{45+4x-x^2}=\frac{2x-1}{(x+5)(x-9)}=\frac{A}{x+5}+\frac{B}{x-9}"

"A(x-9)+B(x+5)=2x-1"

"A+B=2"

"5B-9A=-1"

"5(2-A)-9A=-1"

"A=11\/14,B=17\/14"


"\int\frac{1-2x}{45+4x-x^2}dx=\int(\frac{11}{14(x+5)}+\frac{17}{14(x-9)})dx=\\=\frac{11}{14}log(x+5)+\frac{17}{14}log(x-9)+C"